Michael's Coffee Shop sells a special blend called Mike's Favorite. For the blend, Michael mixes Hazelnut Coffee, which sells for $7.50 per pound, with Cinnamon Coffee, which sells for $6.75 per pound. How much of each flavor must he use to create 10 pounds of Mike's Favorite blend if he sells it for $7.00?
Answer by Nate(3500)

h = the amount of pounds of hazelnut coffee
c = the amount of pounds of cinnamon coffee
*Remember: c = 10 - h
(7.5h + 6.75c)/10 = 7
7.5h + 6.75c = 70
7.5h + 6.75(10 - h) = 70
7.5h + 67.5 - 6.75h = 70
0.75h = 2.5
h = 3 1/3
c = 10 - h = 10 - 3 1/3 = 6 2/3